OpenStudy (anonymous):

i need 2 equations that have 1 solution!!!

OpenStudy (anonymous):

Whats the question

OpenStudy (anonymous):

i need 2 equations that have 1 solution

OpenStudy (nathan917):

1+1=2 4-2=2 :P like this?

OpenStudy (anonymous):

simultanous equations.

OpenStudy (anonymous):

yea but more 9th grade like lol

OpenStudy (anonymous):

2*2=4 2+2=4

OpenStudy (nathan917):

Ill see what i can do @allie_bear22

OpenStudy (anonymous):

okay thnx :)

OpenStudy (anonymous):

Don't just give her the answers

OpenStudy (anonymous):

Explain to her how to do it so she knows how to.

OpenStudy (nathan917):

I said ill see what i can do.. Did i say i would get her the answer? No...

OpenStudy (anonymous):

Am i talking to you. Did you give her a answer. No

OpenStudy (anonymous):

like that @allie_bear22